OpenStudy (anonymous):

Write the equation of the line in standard form: (-3, 6); m = 1/2 (1, 0) & (-5, 3) (-5, -5) & (-1, 3)

OpenStudy (anonymous):

Standard Form is Ax + By = C (1) \[y - 6 = \frac{1}{2} (x+3)\] -->\[x - 2y = -15\] (2) \[y - 0 = m(x-1)\] and\[y - 3 = m(x+5)\] So \[m = \frac{y-3}{x+5}\] Substitute m into the first equation: \[y = \frac{(y-3)(x-1)}{x+5}\] Here you can do simplify by yourself. The result should be: \[x + 2y = 1\] (3) Do as similarly as the 2nd one